40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 4902     Entry time: Tue Jun 28 21:05:05 2011
Author: Jamie 
Type: Update 
Category: SUS 
Subject: SUS control model updated 

I have updated the sus_single_control model, adjusting/cleaning up/simplifying the LSC/POS input signals, and routing new signals to the lockins. Notably one of POS inputs to the part ("lockin_in") was eliminated (see below).

The 6 inputs to the TO_COIL output matrix are now:

LSCPOS + OFFSET + ALT_POS_IN
ASCPIT + OFFSET + SUSPIT + OLPIT
ASCYAW +OFFSET + SUSYAW + OLYAW
SIDE
LOCKIN1
LOCKIN2

The ALT_POS input is used only by the ETMs for the green locking. Just outside of the sus_single_control library part in the ETM models are the green locking controls, consisting of the ETM?_ALS filter bank and the ETM?_GLOCKIN lockin, the outputs from which are summed and fed into the aforementioned ALT_POS input.

As for the SUS lockins (LOCKIN1 and LOCKIN2 in the library model), their input matrix now gets the direct inputs from the OSEMS (before filtering) and the outputs to the coils, after all filtering. These will aid in doing binary output switching tests.

All suspension models (c1sus, c1scx, c1scy) have been rebuild and restarted so that they reflect these changes.

ELOG V3.1.3-